OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME WITH THIS PROPORTIONALITY QUESTIONS!!!?!? I have two sets of numbers: The question asked: Determine whether or not the sugar content is roughly proportional to the caloric value of the beverages. Mass percent solution Calories/mL A 2.8 % 0.2113 Cal/mL B 3.9 % 0 Cal/mL C 8.7 % 0.5072 Cal/mL

OpenStudy (mertsj):

Divide and see if you get roughly the same number.

OpenStudy (mertsj):

And you don't

OpenStudy (anonymous):

I am not sure what I am supposed to divide. I understand proportionality at all.

OpenStudy (mertsj):

y=kx is the form of direct proportionality. So y/x = a constant. So is you divide your y by x, and there is a proportional relationship, you should get close to the same number.

OpenStudy (anonymous):

So the first one I would do this? : 2.8/0.2113 = 13.25

OpenStudy (mertsj):

I would change the percent to a decimal but it doesn't really matter if they are all percents.

OpenStudy (anonymous):

okay i will do that...i get 0.028/0.2113 = .1325 ...Now I don't understand what my result means this is my trouble... what does .1325 mean and how do i know if its propotional to the caloric value.

OpenStudy (mertsj):

Remember I told you that when you divide you should get roughly the same number for all the ordered pairs. And if you don't that means it is not proportional. And that is what you have to answer.

OpenStudy (mertsj):

Determine whether or not the sugar content is roughly proportional to the caloric value of the beverages.

OpenStudy (anonymous):

I think I understand...so i did the calculations and i got for A=0.1325 B= undefined C= 0.1715 Now so does this mean it is no proportional because 0.1325 is not close to 0.2113 and 0.1715 is not close to 0.5072 ?

OpenStudy (anonymous):

Hello are you still there??